# ShareLoRA: Parameter Efficient and Robust Large Language Model Fine-tuning via Shared Low-Rank Adaptation

### v1

@@ 5.3 LLaMA on MMLU
Table 4: Selected the optimal checkpoint based on performance in the five-shot MMLU and evaluated using a zero-shot on MMLU, ARC Challenge, and Hellaswarg, along with a five-shot on GSM8K using the lm-eval-harness leaderboard Gao et al. (2023).
5.4 Zero Shot of ShareA
The effectiveness of ShareA in enhancing generative capabilities is evaluated using both zero-shot and five-shot settings on the lm-eval-harness leaderboard Gao et al. (2023), focusing on tasks like MMLU, ARC Challenge, Hellaswarg, and GSM8K. Results highlight ShareA’s strength in zero-shot learning across various LoRA-configured tasks. ShareA particularly improving performance on domain-specific tasks such as GSM8K that involve mathematical reasoning. This demonstrates ShareA’s robust adaptability and superior performance compared to other models, including the LLaMA 7B, which, despite its strong performance in MMLU as discussed in section 5.3, shows limited adaptability in varied tasks like ARC (c) and GSM8K. Overall, ShareA’s consistency across different domains underscores its effectiveness.
5.5 Quantized ShareLoRA
Method # Params MMLU (5) Method # Params MMLU (5)
LLaMA 7B (QLoRA)* 79.9M 38.8 LLaMA 13B (QLoRA)* 125.2M 47.8
LLaMA 7B (QLoRA)* 79.9M 39.96 LLaMA 13B (QLoRA)* 125.2M 47.29
LLaMA 7B (QLoRA) 79.9M $40.63\pm 0.9$ LLaMA 13B (QLoRA) 125.2M $47.13\pm 0.9$
LLaMA 7B (QShareAqkv) 67.7M $40.63\pm 0.5$ LLaMA 13B (QShareAqkv) 106.0M $\mathbf{47.36\pm 0.7}$
LLaMA 7B (QShareA) 44.6M $\mathbf{41.11\pm 0.2}$ LLaMA 13B (QShareA) 69.5M $47.17\pm 0.8$
Table 5: The performance comparison of LLaMA 7B and 13B with QLoRA and QShareA under the same configuration of Dettmers et al. (2023), * is similar experiment results collected from prior work Xu et al. (2023)
The detailed experiments conducted on training QLoRA for Quantized LLaMA models demonstrate that the QShareA method exhibits better performance compared to QLoRA in general, as shown in the Table5. Despite a reduction in the number of training parameters, both QShareA and QShareAqkv maintain robust and stable in the performance.
Even though, the original weight is quantized and the number of training parameter is further reduced, the performance is not compromised for both QShareA and QShareAqkv. It reveals that the quantization strategies effectively combined with our shared approach without sacrificing output quality.
6 Analysis
6.1 Sharing Attention QKV or Sharing All
@@
The distinction between sharing the self-attention mechanism and all linear modules exists on MLP components like gates and up/down projections, which are suitable for LoRA techniques despite being non-square matrices. This leads to a discrepancy in trainable parameters between LoRA’s A and B. The strategic choice involves deciding whether to uniformly share weights across all layers (ShareA) or selectively share them, such as only for the down projection (ShareAB) while maintaining unique weights for other components like the up projection and gates. Preliminary results in Appendix Figure 4 suggest that selective sharing, particularly of the QKV matrices in Shareqkv, provides an effective balance by aligning closely with both ShareA and LoRA , potentially mitigating overfitting risks.
6.2 Singular Value Decomposition across Layers
@@
As shown in the Figure 6 in Appendix, we apply Singular Value Decomposition (SVD) to the LLaMA 13B both LoRA and ShareA weights. The singular value distributions for the LLaMA 13B model’s LoRA and ShareA weights reveals distinct patterns in their decay rates across layers. The LoRA weights exhibit a sharp decrease in singular values, indicating a concentration of information in a few dominant components, which might lead to specialization and potential overfitting. In contrast, the ShareA weights show a smoother and more gradual decrease, suggesting a more balanced distribution of information among components. This balanced distribution likely enhances the ShareA model’s adaptability and generalization capability across different tasks.
@@ 7 Conclusion
In this paper, we introduce ShareLoRA, a modification of the LoRA architecture that shares either the up or down projection across different layers. The ShareA variant significantly reduces the number of trainable parameters by about half relative to the original LoRA and shows improved performance on fully converged datasets. Through extensive experimentation with NLU, NLG, and zero-shot tasks on models varying from millions to billions of parameters, ShareA provides an optimal balance between computational efficiency and robust performance. By sharing all linear components or focusing solely on self-attention mechanisms, ShareA potentially reduces overfitting risks while maintaining high adaptability and effectiveness across various domains.
@@ Appendix A Hyperparameters
Figure 2: Peak Memory Consumption required for training LLaMA 13B
@@
The hyperparameter setting for RoBERTa is in Table 6 and for LLaMA are in Table 7 and 8.
The number of trainable parameters in Table 5, should remain consistent between QLoRA and LoRA for LLaMA 7B and 13B in Table 3, as both models utilize BFloat16. However, the reduced number of trainable parameters is influenced by the implementation described in Dettmers et al. (2023), which reduces the trainable parameters by half when quantizing to 4 bits. This is also reported the same by Xu et al. (2023), and we maintain this parameter count to ensure consistency.
@@ Appendix B LLaMA Performance Analysis
In Figures 3 and 4 , we present the Dev Set performance changes for both LLaMA and LLaMA2 models, ranging from 7B to 13B, to observe the differences in performance over steps. The results demonstrate that ShareA and ShareAqkv configurations offer several advantages over their counterparts, as discussed in Section 6.1.
@@ Appendix C Convergence Analysis
Figure 6: Distribution of Singular Values for LLaMA 13B: SVD Decomposition Analysis of LoRA (left) and ShareA (right) across All Layers.
@@
Parameters Batch size LR Steps Source Length Target Length
7B 16 2e-4 10000 384 128
13B 16 2e-4 10000 384 128

### v2

@@ 4.2 Baselines
Table 4: Performance of LLaMA models trained on the Alpaca General dataset and tested in a zero-shot of MMLU, ARC Challenge, and Hellaswarg, and in a five-shot of GSM8K, using the lm-eval-harness leaderboard Gao et al. (2023). This table demonstrates the model’s cross-domain adaptability in common sense, reasoning, and mathematics after finetuning on the general dataset.
5 Results
Parameter Efficiency and Performance
ShareLoRA demonstrates significant parameter efficiency while maintaining or improving performance across various model sizes and tasks. For large-scale LLaMA models, as shown in Table 3, ShareA reduces trainable parameters by 44% compared to LoRA. Despite this substantial reduction, ShareA achieves comparable or improved MMLU scores, with LLaMA 13B showing an increase from 47.60 to 48.15.
On the E2E NLG Challenge in Table 2, ShareA demonstrates markedly greater efficiency on GPT-2 models: it reduces LoRA’s parameter count by 43% on the Medium model, yet still achieves performance gains. Specifically, GPT-2 Medium’s BLEU improves from 69.5 to 69.7 and its ROUGE-L from 71.51 to 71.63, while GPT-2 Large’s BLEU increases from 69.8 to 70.0.
Notably, while ShareA consistently outperforms LoRA, our experiments show that ShareB and ShareAB generally underperform compared to ShareA. For instance, in the GPT-2 Large model, Table 2 shows ShareB achieves a BLEU score of 69.7 and a ROUGE-L score of 70.94, which are lower than both LoRA and ShareA.
Comparing ShareLoRA to other state-of-the-art PEFT methods, we observe competitive or superior performance. For instance, on the GLUE benchmark using RoBERTa-large, Table 1 shows ShareA achieves an average score of 88.6 on GLUE, compared to 84.9 for Prefix-tuning while using significantly fewer parameters. Even in its most aggressive configuration, ShareAB, with only 0.03M trainable parameters which reduces 96% trainable parameters compared to LoRA, outperforms IA3 which uses 0.18M parameters, achieving an average score of 87.0 on GLUE compared to IA3’s 86.0. Furthermore, under a similar trainable parameter size, ShareA demonstrates better performance than LoRA-FA. For example, ShareA achieves an average GLUE score of 89.3 with 0.4M parameters on RoBERTa-large, surpassing LoRA-FA’s score of 88.5 with the same parameter count.
Model Adaptability
ShareLoRA demonstrates superior adaptability across a diverse range of tasks and model sizes. In experiments with RoBERTa-base model on the GLUE benchmark shown in Table 1, ShareA exhibits particular strength on smaller datasets that are typically prone to overfitting. Specifically, on tasks such as MRPC, CoLA, and RTE, ShareA achieves performance gains of 0.2% to 0.5%. These improvements are especially noteworthy given that these datasets have generally reached full convergence under standard training configurations Hu et al. (2021), suggesting ShareLoRA’s ability to extract additional performance even in challenging scenarios.
Method MATH MMLU MMLU Pro
LLaMA3 8B (LoRA) 12.46 65.93 31.80
LLaMA3 8B (ShareA) 13.24 66.35 32.55
LLaMA3.1 8B (LoRA) 15.36 65.59 32.86
LLaMA3.1 8B (ShareA) 15.10 65.78 33.69
Table 5: Performance of LLaMA3 trained on the MATH dataset (Hendrycks et al., 2021b) and evaluated in a zero-shot of MATH and in five-shot of MMLU and MMLU-Pro. It highlights the model’s ability to maintain cross-domain adaptability in common sense and reasoning domains after finetuning on mathematics.
LLaMA3 8B Phase 1 ALPACA Phase 2 GSM8K Phase 3 CodeALPACA Phase 4 ALPACA
Tasks MMLU MMLU Pro GSM8K HumanEval MMLU MMLU Pro
LoRAqv 65.44 33.15 55.64 38.41 65.32 33.14
ShareLoRAqv 65.94 33.85 56.78 39.02 65.30 33.36
Table 6: Continual Adaption across multiple tasks, starting with Alpaca, followed by GSM8K, then CodeAlpaca, and finally revisiting Alpaca. At each stage, we evaluate the effectiveness of continual adaptation by leveraging the best checkpoints from the preceding task, comparing both LoRA and ShareLoRA for LLaMA3 8B.
Figure 2: Memory Consumption of LLaMA3 70B with QLoRA and QLoRA-shareA (QShareA).
ShareA further showcases enhanced transfer learning capabilities. When fine-tuning on adaptive tasks like MRPC, RTE, and STS-B using the best MNLI checkpoint, ShareA consistently performs on par with or outperforms LoRA. Notably, ShareA outperforms other PEFT methods in this transfer learning scenario as well. For instance, on the RTE task, ShareA, with 0.16M parameters for RoBERTa-base, achieves a score of 87.1, significantly surpassing Prefix-tuning’s 54.51 as shown in Table 1. ShareA also demonstrates superior performance when compared to methods with similar trainable parameter sizes, such as BitFit with 0.1M parameters and LoRA-FA with 0.15M parameters. This highlights ShareA’s efficiency in parameter utilization and its ability to extract better performance from a given parameter budget, particularly in transfer learning scenarios.
Robustness Across Domains
ShareLoRA shows strong robustness and adaptability across both diverse task domains and varying model sizes. As presented in Tables 3 and 4, ShareLoRA consistently surpasses LoRA in zero-shot and few-shot learning scenarios across multiple evaluation benchmarks.
On the LLaMA2-7B model, ShareLoRA improves MMLU accuracy by 0.7%, while on the LLaMA2-13B model, it achieves a 0.5% gain. Beyond MMLU, ShareLoRA delivers average performance gains of 1.8% and 1.3% on LLaMA2-7B and LLaMA2-14B models, respectively, with accuracy improvements ranging from 0.5% to 2.5% across various tasks. These results collectively underscore ShareLoRA’s effectiveness in enhancing model generalization and transferability across both small and large-scale language models.
Continual Adaptation
To assess the robustness and knowledge retention during continual fine-tuning, we deploy the LLaMA3 and LLaMA3.1 models on the MATH dataset. We then evaluate their performance in mathematics and across other domains, such as MMLU and MMLU-Pro, to compare how well these models preserve knowledge, as shown in Table 5. Our findings indicate that both ShareLoRA and LoRA deliver matched performances for directly fine-tuned domains. However, when adapting these fine-tuned models to other evaluation benchmarks, ShareLoRA demonstrates greater robustness, outperforming LoRA. Specifically, on MMLU-Pro, ShareLoRA outperforms LoRA by 0.86% on LLaMA3.1 and 0.75% on LLaMA3.
We also investigate continual fine-tuning across multiple tasks—starting from Alpaca, followed by GSM8K, then CodeAlpaca, and finally returning to Alpaca in Table 6. ShareLoRA consistently outperforms LoRA in this setting, with observed gains of 0.5% on MMLU and MMLU-Pro, 1.2% on GSM8K, and 0.6% on HumanEval, highlighting its robustness in multi-task continual learning.
6 Analysis and Discussion
Relative Importance of LoRA Components
Our experimental findings demonstrate that both LoRA and ShareA consistently outperform ShareB in a variety of classification and generative tasks, across most metrics. Within the LoRA framework, the up-projection matrix B plays a pivotal role by significantly enhancing the dimensionality of the low-rank representation. Consequently, it is both practical and justifiable to share the less critical module, LoRA A, while retaining the integrity of B. However, sharing both matrices A and B simultaneously tends to compromise too much critical information. Particularly in generative tasks, opting to share component A instead of B within the ShareLoRA framework is strategically beneficial, as seen in Table 2. This is because expanding the intermediate dimension proves more crucial and challenging than compressing high-dimensional features in complex generative scenarios.
Sharing Attention QKV vs. Sharing All
@@ 6 Analysis and Discussion
The distinction between sharing the self-attention mechanism and all linear modules exists on MLP components like gates and up/down projections. This leads to a discrepancy in trainable parameters between LoRA’s A and B. The strategic choice involves deciding whether to uniformly share weights across all layers (ShareA) or selectively share them, such as only for the down projection (ShareAB) while maintaining unique weights for other components like the up projection and gates. Preliminary results in Appendix Figure 5 suggest that selective sharing, particularly of the QKV matrices in Shareqkv, provides an effective balance by aligning closely with both ShareA and LoRA , potentially mitigating overfitting risks.
Memory Footprint
In the context of smaller models like RoBERTa and GPT-2, ShareA yields minimal parameter savings, which is negligible given modern GPU capacities. However, with larger models like LLaMA, ShareA demonstrates more substantial reductions. Specifically, the LLaMA 7B and 13B models cut down approximately 60 million and 110 million trainable parameters, when compared to the LoRA. This leads to substantial efficiency gains, reducing both computational footprint and disk storage needs.
As depicted in Figure 2 and Figure 7, in the Llama3 70B model, the ShareA adaptation achieves a 6.3GB reduction in memory footprint under the quantization configuration. Meanwhile, in the Llama2 13B model with the LoRA configuration, ShareA manages to reduce the memory footprint by 3.8GB and enhances training speed by approximately 3%. The confidence intervals in Table 3 illustrate that ShareA not only improves performance but also increases robustness over standard LoRA, underscoring the practical advantages of ShareLoRA in LLMs.
SVD Analysis of LoRA and ShareA Weights
@@
We conducted a Singular Value Decomposition (SVD) analysis on the weights of LLaMA 13B for both LoRA and ShareA, as shown in Figure 3 in the Appendix. The results reveal distinct patterns in their singular value distributions across layers. LoRA weights exhibit a sharp decrease in singular values, indicating a concentration of information in a few dominant components. This could lead to specialization but might also increase the risk of overfitting. In contrast, ShareA weights show a smoother, more gradual decrease in singular values, suggesting a more balanced distribution of information among components. This balanced distribution contributes to ShareA’s enhanced adaptability and generalization capability across different tasks.
These findings provide insight into why ShareA may offer improved robustness and continue training performance compared to LoRA. The more uniform singular values distribution in ShareA suggests that it captures richer features, leading to better generalization across various domains.
@@ 7 Conclusion
In this paper, we introduce ShareLoRA, an optimization of the LoRA architecture that shares either the up or down projection across different layers. ShareLoRA significantly reduces the number of trainable parameters by at least half relative to the original LoRA and shows improved performance on fully converged datasets. Through extensive experimentation with NLU, NLG, and zero-shot tasks on models of varying scales, ShareLoRA demonstrates a strong balance between computational efficiency and robust performance. It consistently maintains high adaptability, strong robustness, and effective continual learning capabilities across diverse tasks and architectures.
@@ Appendix A Hyperparameters
The hyperparameter setting for RoBERTa is in Table 8 and for LLaMA are in Table 10 and 11. The number of trainable parameters in Table 7, should remain consistent between QLoRA and LoRA for LLaMA 7B and 13B in Table 3, as both models utilize BFloat16. However, the reduced number of trainable parameters is influenced by the implementation described in Dettmers et al. (2023), which reduces the trainable parameters by half when quantizing to 4 bits. This is also reported the same by Xu et al. (2023), and we maintain this parameter count to ensure consistency.
@@ Appendix B LLaMA Performance Analysis
In Figures 4 and 5, we present the Dev Set performance changes for both LLaMA and LLaMA2 models, ranging from 7B to 13B, to observe the differences in performance over steps. The results demonstrate that ShareA and ShareAqkv configurations offer several advantages over their counterparts, as discussed in Section 6.
@@
The robustness of ShareLoRA extends to quantized models. Table 7 shows that QShareA (QLoRA with ShareA) maintains strong performance even with substantial parameter reduction. In the case of LLaMA 7B, QShareA achieves an MMLU score of 41.11, surpassing QLoRA’s score of 40.63. This trend continues with larger models: for LLaMA 13B, QShareA slightly outperforms QLoRA with scores of 47.17 and 47.13 respectively, while using significantly fewer parameters. These performance gains are consistently observed across different model sizes, including LLaMA2 7B and LLaMA 13B, highlighting ShareLoRA’s broad applicability and scalability.
@@ Appendix C Convergence Analysis
Dataset E2E Challege
Optimizer AdamW
Weight Decay 0.01
Dropout Prob 0.1
Batch Size (per device) 8
# Epochs 5
Warmup Steps 500
Learning Rate Schedule Linear
Label Smooth 0.1
Learning Rate 0.002
Adaptation $r_{q}=r_{v}=4$
LoRA $\alpha$ 32
Beam Size. 10
Length Penalty 0.9
no repeat ngram size 4
Table 9: Configuration and training details for GPT-2 LoRA on E2E Challenge
Model Parameters Batch size LR Steps Source Length Target Length LoRA r LoRA $\alpha$
LLaMA1 & 2 7B 16 2e-4 10000 384 128 64 16
LLaMA1 & 2 13B 16 2e-4 10000 384 128 64 16
LLaMA3 & 3.1 8B 16 2e-5 5000 384 128 64 16
@@
Figure 7: Memory Consumption required for LLaMA2 13B.
Appendix D Memory Footprint
We utilizes float32 for QLoRA modules to enhance the performance of quantized models, while bfloat16 is employed for LoRA fine-tuning. We employ the standard AdamW optimizer with a batch size of 1, a sequence length of 512, and do not use gradient checkpointing.
The chart in Figure  7 depicts memory usage across four configurations of the Llama2 13B model: LoRA, LoRA-Shared A, QLoRA, and QLoRA-Shared A, highlighting the impact of model scaling and adaptations on resource needs. It shows a memory reduction of 3.8 GB when using LoRA-Shared A compared to the LoRA configuration, and a further savings of 2.1 GB with QLoRA-Shared A compared to QLoRA. LoRA-Shared operates independently from QLoRA strategies, thereby reducing memory usage further without interfering with LoRA or QLoRA configurations.